OpenStudy (erinweeks):

Havent done this in a while.. Write a formula for the general term (the nth term) of the geometric sequence.

OpenStudy (erinweeks):

\[\frac{ 1 }{ 2 }, \frac{ 1 }{ -10 }, \frac{ 1 }{ 50 }, \frac{ 1 }{ -250 }\]

OpenStudy (erinweeks):

i so totally forgot how to do this !

OpenStudy (jack1):

bear with me... its been a while for me to geom seq \[a _{n} = a \times r ^{n-1}\]??? i think?

OpenStudy (erinweeks):

i dont even know lol .

OpenStudy (jack1):

a = ur initial r = ur ratio or constant or whatever u call it n = the term in the sequence you want

OpenStudy (erinweeks):

n would be -5 ?

OpenStudy (jack1):

1 over -5...? term 2 divided by term 1

OpenStudy (jack1):

so 1/-10 divided by 1/2 = 1/-5

OpenStudy (erinweeks):

or r would be -5

OpenStudy (erinweeks):

okay

OpenStudy (jack1):

r = 1/-5... n = whichever term we want

OpenStudy (jack1):

a = 1/2 ... im pretty sure that's all we need for the equation hey ;D

OpenStudy (erinweeks):

hold on im going to post the answers

OpenStudy (erinweeks):

\[A. a \eta = (\frac{ 1 }{ 2 }) n-1 \frac{ 3 }{ 5 }\]

OpenStudy (erinweeks):

B.